Cassini Significant Events 04/18/07 - 04/24/07

April 26, 2007

(Source: Cassini Project)


The most recent spacecraft telemetry was acquired on Tuesday, April 24, from
the Goldstone tracking complex. The Cassini spacecraft is in an excellent
state of health and all subsystems are operating normally. Information on
the present position and speed of the Cassini spacecraft may be found on the
"Present Position" web page located at
http://saturn.jpl.nasa.gov/operations/present-position.cfm.

Wednesday, April 18 (DOY 108):

Orbit Trim Maneuver (OTM) #105 was performed today. This is the apoapsis
maneuver setting up for the Titan 29 encounter on April 26. The main engine
burn began at 9:45 PM PST. Telemetry immediately after the maneuver showed
the burn duration was 21.9 seconds, giving a delta-V of 3.51 m/s. All
subsystems reported nominal performance after the maneuver.

Friday, April 20 (DOY 110):

An encounter strategy meeting was held today to cover the period between
April 26 and May 12, Titan flybys T29 and T30, and maneuvers 107-109.

Monday, April 23 (DOY 113):

Orbit Trim Maneuver (OTM) #106 was performed today. This is the approach
maneuver setting up for the Titan 29 encounter on April 26. The reaction
control subsystem (RCS) burn began at 9:30 PM PST. Telemetry immediately
after the maneuver showed the burn duration was 9.0 seconds, giving a
delta-V of 0.017 m/s. All subsystems reported nominal performance after the
maneuver.

This was the first OTM where the team adjusted the targeted time of arrival
in order to make the OTM large enough to perform. As a policy, we had set
the minimum delta-V limit for RCS burns at about 10 mm/sec. Without the
arrival time adjustment, this OTM was around 8 mm/sec., which is below the
minimum limit. The team was further constrained in that this OTM was not a
candidate for cancellation, as the downstream delta-V cost of more than 2
m/sec was too high. This OTM was also particularly challenging as the team
tried to find a design that minimized low-rpm regions for the reaction
wheels.

The Science Operations Plan update process for the S33 sequence kicked off
today. The process will run for about five weeks and then will be handed
off to the sequence leads for final development.

Tuesday, April 24 (DOY 114):

Today there were non-targeted flybys of Dione and Telesto. Although there
were no observations planned for Telesto, the Imaging Science Subsystem
acquired a two-by-two narrow angle camera mosaic of Dione, referred to as a
regional map.

The Cassini-Huygens Analysis and Results of the Mission (CHARM)
teleconference for April was held today. The topic, "Enceladus: Cassini
finds another Active World" was presented by a member of the Composite
Infrared Spectrometer instrument team.

The sequence for the second Radio and Plasma Wave Science Whistler search
was uplinked to the spacecraft last Friday and executed on board today. The
science teams are in the process of analyzing the data.